OpenStudy (chris215):

dy/dx for x^2 + y^2 = 2xy.

OpenStudy (freckles):

find y'?

OpenStudy (chris215):

i got 1

OpenStudy (freckles):

you can make this easy by -2xy on both sides... \[x^2-2xy+y^2=0 \\ (x-y)^2=0 \\ x-y=0 \\ y=x\]

OpenStudy (freckles):

then find y'

OpenStudy (freckles):

and yes y'=1 since y=x

OpenStudy (chris215):

thank you so much!

OpenStudy (freckles):

np you could find that y'=1 doing what I call the long way around :p

OpenStudy (freckles):

which is the way I bet you went about it